Due to annihilation of electron- positron of same kinetic energy 0.95 MeV, a photon is produced which canalso be produced by an electron of energy E, the possible value of E is ( mass of electron = 9.1 x 10-31 kg, e = 1.6 x 10-19C )
1
Expert's answer
2021-08-15T18:21:02-0400

We know that

"E=K+m_0c^2"

"K=E-m_0c^2"

Put value


"E=(0.95+9.1\times10^{-31}\times(3\times10^8)^2\times931)MeV"


1amu=931 MeV



"E=(0.95+7.62\times10^{-11})MeV"

"E=0.95MeV"